# 0.5pt (eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version

### v2

@@ 5.2 Quantitative Results
5.3 Qualitative Results
Figure 4 shows examples of synthesized images with PyNeRF [52] and Instant-NGP [34] in the sparser setting, with and without our method. As seen in the quantitative results (Section 5.2), the output quality of the proposed Re-Nerfing improves significantly upon the baselines, especially in terms of the reduction of floating artifacts ("floaters"). In particular, the addition of our method sharpens the scene and better preserves the small details, such as the table cloth in the bonsai scene, or the books and picture frame in the room scene. Remarkably, the proposed method is able to remove dominant artifacts in the renderings by Instant-NGP [34] and smaller ones from PyNeRF [52], despite using the same original views, architecture, and hyperparameters as the baselines. The augmented views help disambiguate the scene geometry, allowing both models to deliver better results.
In Figure 1, we show additional qualitative results, including examples from 3D Gaussian Splatting [26]. The improvements over the base 3DGS are remarkable. Our augmentation method brings significant improvements, resulting in sharper renderings across the board, with 3D Gaussians better fitting the scene geometry.
The efficacy of our method on both implicit (NeRF-based) and explicit scene representations is particularly notable as it is achieved without using extra data, additional models, or supervision, only using already available information.
Figure 4: Qualitative results on cropped images from the test set of the mip-NeRF 360 dataset. The models were trained on the sparser setting. The proposed Re-Nerfing is applied on PyNeRF [52] and Instant-NGP [34].
The supplementary material includes additional details and results, such as the performance on each scene, the ability of Re-Nerfing to improve on specific views, and extra qualitative examples in a video.
Limitations Since our method relies on the first stage to provide reasonable renderings from a good estimate of the underlying scene geometry, it is challenging to enhance poorly rendered scenes. In extremely sparse settings where the rendering quality is lower, combining our approach with other methods [51, 15] would be an interesting direction. The proposed method also relies on the uncertainty estimates to discard artifacts and low-quality regions. Therefore, better uncertainty estimates would be beneficial in the second and later stages. Uncertainty calibration could help refine the estimates while preserving the signal from higher-quality areas.
6 Conclusion
This work enhances NeRFs and 3DGSs by leveraging their own novel view synthesis capabilities. The introduced Re-Nerfing is a general approach that synthesizes novel views to improve the scene coverage and augment the training data for a new model. To provide a valuable training signal for the second model, Re-Nerfing computes the NeRF’s uncertainty for each augmented view to discard the uncertain regions and reduce artifacts. As shown with various pipelines, this simple method yields improvements in both rendering quality and convergence speed of sparser and denser scenes, mitigating the need for extensive captures of the scene.
Appendix 0.A Supplementary Material
In this supplementary material, we complement the main paper with additional results in greater detail. Specifically, in Section 0.A.1 we show how our method can optimize for specific viewpoints, in Section 0.A.2, we show the results for each scene separately, and in Section 0.A.3 we discuss qualitative results from the attached supplementary video.
0.A.1 Optimizing for Specific Views
sparser views
Method PSNR SSIM LPIPS
PyNeRF [52] 22.42 0.632 0.218
adding 4 randomly interpolated synthetic views (1.08x) 22.58 0.653 0.205
adding 4 test synthetic views (1.08x) 22.64 0.663 0.196
Table 5: View optimization strategies on the sparser setting of the stump scene of the mip-NeRF 360 [3] dataset (50 views). All models are based on PyNeRF [52].
In Table 5, we show how our method offers the opportunity of optimizing for specific views. Toward this end, we train PyNeRF [52] augmenting with 4 randomly interpolated views synthesized by the baseline and compare it with another PyNeRF model trained augmenting with 4 test views synthesized by the same baseline. Both models are evaluated on the viewpoint of the 4 test views. While adding the randomly interpolated views benefits the output quality, specifically synthesizing the viewpoints of interest (i.e., the test views in this case) shows a higher margin on all metrics. In practice, this means that Re-Nerfing allows to improve on specific viewpoints of interest, which can be beneficial when the views available lead to a sub-optimal quality from specific areas. In such setting, Re-Nerfing can be used to densify the viewpoints around the area lacking in terms of quality, and ultimately improve the novel view synthesis.
0.A.2 Evaluation by Scene
In Tables from 6 to 12, we report the performance for each scene for the various pipelines on which we applied Re-Nerfing, namely Instant-NGP [34], PyNeRF [52], and 3D Gaussian Splatting [26]. The detailed results showcase the benefits of our method indoor and outdoor, for sparser and denser view settings, and for both implicit and explicit methods.
0.A.3 Additional Qualitative Results
@@
We point the reader to the attached supplementary video https://youtu.be/8mo3s1XAwHQ for more qualitative results. The video interpolates across the test views. Remarkably, when paired with the proposed Re-Nerfing  3DGS [26] reaches significantly higher rendering quality, even in the sparser setting showcased in the video for the counter scene (30 views). Thanks to our method, the scene geometry is substantially better captured by the Gaussians, eliminating major structural artifacts. As shown in Table 9, the garden scene is easier to render. In this case, our method helps with the scene geometry, especially visible in the rendered depths, as it eliminates the holes in the ground.
@@
+ Re-Nerfing [ours]* 21.08 0.686 0.241 27.70 0.879 0.127
Table 6: Results on the counter scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on Instant-NGP [34], PyNeRF [52], and 3D Gaussian Splatting (3DGS) [26]. *: no uncertainty estimates were used.
@@
+ Re-Nerfing [ours]* 18.36 0.324 0.377 25.39 0.774 0.149
Table 7: Results on the bicycle scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on Instant-NGP [34], PyNeRF [52], and 3D Gaussian Splatting (3DGS) [26]. *: no uncertainty estimates were used.
@@
+ Re-Nerfing [ours]* 19.97 0.433 0.306 24.32 0.691 0.172
Table 8: Results on the stump scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on PyNeRF [52] and 3D Gaussian Splatting (3DGS) [26]. Instant-NGP [34] could not deal with this scene in the sparser setting (PSNR of 15.55), so it was not applied here. *: no uncertainty estimates were used.
@@
+ Re-Nerfing [ours]* 23.18 0.730 0.106 27.93 0.900 0.048
Table 9: Results on the garden scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on Instant-NGP [34], PyNeRF [52], and 3D Gaussian Splatting (3DGS) [26]. *: no uncertainty estimates were used.
@@
+ Re-Nerfing [ours]* 23.09 0.705 0.217 30.61 0.931 0.087
Table 10: Results on the bonsai scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on Instant-NGP [34], PyNeRF [52], and 3D Gaussian Splatting (3DGS) [26]. *: no uncertainty estimates were used.
@@
+ Re-Nerfing [ours]* 22.56 0.797 0.130 30.21 0.928 0.064
Table 11: Results on the kitchen scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on Instant-NGP [34], PyNeRF [52], and 3D Gaussian Splatting (3DGS) [26]. *: no uncertainty estimates were used.
@@
+ Re-Nerfing [ours]* 20.89 0.715 0.271 30.02 0.907 0.128
Table 12: Results on the room scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on Instant-NGP [34], PyNeRF [52], and 3D Gaussian Splatting (3DGS) [26]. *: no uncertainty estimates were used.

### v3

@@ A.3 Dense-View Settings
A.4 Additional Ablation Study
In Table 8, we report an ablation study for both PyNeRF [59] and Instant-NGP [38] in the 30-view setting. Training again with the synthetic views produced by the baseline (with $N=7$) slightly degrades the performance for all metrics for PyNeRF, while it improves for Instant-NGP. The degradation for PyNeRF is due to the lower quality signal of 6x more views than the reliable signal of the original views. Specifically, the model learns the artifacts of the augmented views and underperforms. Instead, for Instant-NGP, the quality of the augmented views was higher, leading to improvements. Nevertheless, removing the images from the sampling pool once the scene structure has been mainly learned (stop, 3rd row) prevents both models from overfitting to the artifacts and leads to better results on all metrics, with large improvements for PyNeRF. Furthermore, masking out the uncertain regions of such synthetic views improves even further across the board for both methods. This shows once again how synthetic views of the same scene are beneficial aid the novel view synthesis task. In particular, the results confirm the importance of increasing the scene coverage and that of having a good training signal, testifying the need for the uncertainty-based masks and our augmentation strategy.
A.5 Additional Qualitative Results
PyNeRF In Figure 4, we showcase the impact of our method on PyNeRF [59]. As for the other methods seen in the main paper, our Re-Nerfing reduces artifacts and leads to a better solution, e.g., by removing floaters. Examples are the floating leaves of the bonsai (first and last rows), or the black blob (second row).
@@ A.5 Additional Qualitative Results
Video We refer to the attached supplementary video for more qualitative results. The video interpolates across the test views for models trained in sparse settings. Remarkably, when paired with the proposed Re-Nerfing  3DGS [28] reaches significantly higher rendering quality in the counter scene (30 views). Thanks to our method, the scene geometry is substantially better captured by the Gaussians, eliminating major structural artifacts. As shown in Table 20, the garden scene is easier to render. In this case, our method helps with the scene geometry, especially visible in the rendered depths, as it eliminates the holes in the ground.
A.6 Evaluation by Scene
LLFF In Tables from 9 to 16, we report the performance for each scene in the 3-view setting of LLFF [36] for RegNeRF [40] and Re-Nerfing. The detailed results showcase that the proposed method always improves upon RegNeRF, while the performance sometimes saturates before 3 rounds.
mip-NeRF 360 In Tables from 17 to 23, we report the performance for each scene for the various pipelines on which we applied Re-Nerfing, namely Instant-NGP [38], PyNeRF [59], and 3D Gaussian Splatting [28]. The detailed results showcase the benefits of our method indoor and outdoor, for sparser and denser view settings, and for both implicit and explicit methods.
fortress, 3 views PSNR SSIM LPIPS
RegNeRF [40] 23.32 0.743 0.065
+ Re-Nerfing, 1 round [ours] 23.69 0.789 0.054
+ extra round, 2nd [ours] 23.94 0.814 0.047
+ extra round, 3rd [ours] 24.34 0.829 0.046
Table 9: Results on the fortress scene of the LLFF dataset [36] for the 3-view settings. The proposed Re-Nerfing is applied on RegNeRF [40] and then to itself iteratively (rounds).
horns, 3 views PSNR SSIM LPIPS
RegNeRF [40] 15.65 0.610 0.200
+ Re-Nerfing, 1 round [ours] 15.75 0.654 0.179
+ extra round, 2nd [ours] 15.80 0.655 0.168
+ extra round, 3rd [ours] 16.68 0.670 0.161
Table 10: Results on the horns scene of the LLFF dataset [36] for the 3-view settings. The proposed Re-Nerfing is applied on RegNeRF [40] and then to itself iteratively (rounds).
fern, 3 views PSNR SSIM LPIPS
RegNeRF [40] 19.87 0.697 0.202
+ Re-Nerfing, 1 round [ours] 19.96 0.709 0.191
+ extra round, 2nd [ours] 19.96 0.714 0.180
+ extra round, 3rd [ours] 19.95 0.718 0.180
Table 11: Results on the fern scene of the LLFF dataset [36] for the 3-view settings. The proposed Re-Nerfing is applied on RegNeRF [40] and then to itself iteratively (rounds).
flower, 3 views PSNR SSIM LPIPS
RegNeRF [40] 19.93 0.688 0.156
+ Re-Nerfing, 1 round [ours] 20.28 0.711 0.143
+ extra round, 2nd [ours] 20.46 0.720 0.140
+ extra round, 3rd [ours] 20.04 0.695 0.154
Table 12: Results on the flower scene of the LLFF dataset [36] for the 3-view settings. The proposed Re-Nerfing is applied on RegNeRF [40] and then to itself iteratively (rounds).
orchids, 3 views PSNR SSIM LPIPS
RegNeRF [40] 15.56 0.502 0.177
+ Re-Nerfing, 1 round [ours] 15.77 0.526 0.162
+ extra round, 2nd [ours] 15.86 0.540 0.155
+ extra round, 3rd [ours] 15.90 0.547 0.150
Table 13: Results on the orchids scene of the LLFF dataset [36] for the 3-view settings. The proposed Re-Nerfing is applied on RegNeRF [40] and then to itself iteratively (rounds).
leaves, 3 views PSNR SSIM LPIPS
RegNeRF [40] 16.60 0.612 0.132
+ Re-Nerfing, 1 round [ours] 16.92 0.654 0.116
+ extra round, 2nd [ours] 17.07 0.670 0.107
+ extra round, 3rd [ours] 17.14 0.677 0.103
Table 14: Results on the leaves scene of the LLFF dataset [36] for the 3-view settings. The proposed Re-Nerfing is applied on RegNeRF [40] and then to itself iteratively (rounds).
room, 3 views PSNR SSIM LPIPS
RegNeRF [40] 21.53 0.861 0.116
+ Re-Nerfing, 1 round [ours] 21.84 0.873 0.105
+ extra round, 2nd [ours] 21.96 0.878 0.098
+ extra round, 3rd [ours] 21.47 0.871 0.107
Table 15: Results on the room scene of the LLFF dataset [36] for the 3-view settings. The proposed Re-Nerfing is applied on RegNeRF [40] and then to itself iteratively (rounds).
t-rex, 3 views PSNR SSIM LPIPS
RegNeRF [40] 20.16 0.766 0.133
+ Re-Nerfing, 1 round [ours] 20.55 0.788 0.117
+ extra round, 2nd [ours] 20.70 0.797 0.111
+ extra round, 3rd [ours] 20.63 0.796 0.113
Table 16: Results on the t-rex scene of the LLFF dataset [36] for the 3-view settings. The proposed Re-Nerfing is applied on RegNeRF [40] and then to itself iteratively (rounds).
@@ A.6 Evaluation by Scene
+ Re-Nerfing [ours] 21.08 0.686 0.241 27.70 0.879 0.127
Table 17: Results on the counter scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on Instant-NGP [38], PyNeRF [59], and 3DGS [28]. Uncertainty estimates are unavailable for 3DGS.
@@
+ Re-Nerfing [ours] 18.36 0.324 0.377 25.39 0.774 0.149
Table 18: Results on the bicycle scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on Instant-NGP [38], PyNeRF [59], and 3DGS [28]. Uncertainty estimates are unavailable for 3DGS.
@@
+ Re-Nerfing [ours] 19.97 0.433 0.306 24.32 0.691 0.172
Table 19: Results on the stump scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on PyNeRF [59] and 3DGS [28]. Instant-NGP [38] could not deal with this scene in the sparser setting (PSNR of 15.55), so it was not applied here. Uncertainty estimates are unavailable for 3DGS.
@@
+ Re-Nerfing [ours] 23.18 0.730 0.106 27.93 0.900 0.048
Table 20: Results on the garden scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on Instant-NGP [38], PyNeRF [59], and 3DGS [28]. Uncertainty estimates are unavailable for 3DGS.
@@
+ Re-Nerfing [ours] 23.09 0.705 0.217 30.61 0.931 0.087
Table 21: Results on the bonsai scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on Instant-NGP [38], PyNeRF [59], and 3DGS [28]. Uncertainty estimates are unavailable for 3DGS.
@@
+ Re-Nerfing [ours] 22.56 0.797 0.130 30.21 0.928 0.064
Table 22: Results on the kitchen scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on Instant-NGP [38], PyNeRF [59], and 3DGS [28]. Uncertainty estimates are unavailable for 3DGS.
@@
+ Re-Nerfing [ours] 20.89 0.715 0.271 30.02 0.907 0.128
Table 23: Results on the room scene of the mip-NeRF 360 dataset [3], for sparser and denser settings. The proposed Re-Nerfing is applied on Instant-NGP [38], PyNeRF [59], and 3DGS [28]. Uncertainty estimates are unavailable for 3DGS.